OpenStudy (anonymous):

factor the trinomial: (3x+2)^2 + 8 (3x+2) + 12

OpenStudy (anonymous):

1) Simplify 2) Find roots of (3x+2)^2 + 8 (3x+2) + 12 = 0 3) Answer is 9(x-1st root)(x-2nd root)

OpenStudy (anonymous):

1st step is not necessary.

OpenStudy (anonymous):

Answer is (3x+4) (3x+8)

OpenStudy (anonymous):

Yes, it can also be given in form 9(x+4/3)(x+8/3)

OpenStudy (anonymous):

The roots are -4/3 and -8/3

OpenStudy (anonymous):

9(x+4/3)(x+8/3) = 3*(x+4/3)*3(x+8/3) = (3x+4) (3x+8)

OpenStudy (zarkon):

(3x+2)^2 + 8 (3x+2) + 12 Let \[u=3x+2\] \[u^2+8u+12=(u+2)(u+6)\] \[=(3x+2+2)(3x+2+6)=(3x+4)(3x+8)\]